# Quantum Metrology via Floquet-Engineered Two-axis Twisting and Turn Dynamics

### v1

@@ Abstract
Optimal measurement with Floquet-engineered anti-TAT-and-turn.— It is crucial to approach the ultimate precision bound (i.e. the QCRB) set by the QFI. For a tiny estimated phase, time-reversal readout may approach the QCRB without single-particle resolved detection [69, 70, 71, 72, 68, 73, 74, 75, 76, 77, 78, 79, 80, 81]. Here, we find that the optimal measurements can be achieved via anti-TAT-and-turn, regarded as the time-reversal process of TAT-and-turn. More importantly, the anti-TAT-and-turn can be realized by Floquet engineering without the need of flipping the sign of nonlinear interaction $\chi$.
@@
With the generated non-Gaussian entangled state, one can encode the estimated phase along the best sensing direction by $\hat{R}_{\vec{n}_{\rm max}}(\phi)=e^{-i\phi\hat{J}_{\vec{n}_{\rm max}}}$. Based upon the Hamiltonian (9), adjusting the frequency of coupling field and the phase of Rabi frequency, $\delta\mapsto-\delta$ and $\alpha=0\mapsto\alpha=\pi/2$, one can obtain an effective anti-TAT-and-turn Hamiltonian $\hat{H}_{\rm{a-TATNT}}^{\alpha=0}=-\hat{H}_{\rm TATNT}^{\alpha=0}$. Thus the state before measurement can be written as
@@
where $\lvert\psi_{i}\rangle=\lvert\uparrow\rangle^{\otimes N}$, $\hat{H}_{\rm TATNT}^{\alpha=0,\delta}=\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})+\delta_{\rm eff}\hat{J}_{z}$ and $\hat{H}_{\rm{TATNT}}^{\alpha=\pi/2,-\delta}=-\hat{H}_{\rm{TATNT}}^{\alpha=0,\delta}=-\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})-\delta_{\rm eff}\hat{J}_{z}$. Therefore, to find out the optimal measurement, one can optimize the sensitivity over all possible directions through singular value decomposition
@@
where $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ with two $3\times 3$ matrices $\mathcal{M}$ and $\mathcal{Q}$, whose elements are given as $\mathcal{M}_{\mu\nu}=i\langle[\hat{U}^{\dagger}_{1}\hat{J}_{\mu}\hat{U}_{1},\hat{\tilde{J}}_{\nu}]\rangle_{i}$ and $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle_{i}/2-\langle\hat{\tilde{J}}_{\mu}\rangle_{i}\langle\hat{\tilde{J}}_{\nu}\rangle_{i}$ for $\mu,\nu=x,y,z$ with $\hat{\tilde{J}}_{\mu(\nu)}=\hat{U}^{\dagger}_{1}\hat{R}_{\vec{n}_{\rm max}}^{\dagger}(\phi)\hat{U}^{\dagger}_{2}\hat{J}_{\mu(\nu)}\hat{U}_{2}\hat{R}_{\vec{n}_{\rm max}}(\phi)\hat{U}_{1}$ [94, 95]. The optimal measurement direction is $\vec{m}_{\rm max}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with the normalization constant $\mathcal{N}_{c}$ determined by the Cauchy-Schwarz inequality [66]. Thus one can easily rotate the final state and measure the half-population difference $\hat{J}_{z}$ to infer $\phi$. Notably, the above calculations are valid for arbitrary phase $|\phi|\geq 0$, which generalize the result only for the zero phase $\phi=0$ [94, 95]. As shown in Fig. 3 (a), the nonlinear readout with Floquet-engineered TAT-and-turn achieves better measurement precision than those with TAT, OAT-and-turn and OAT. The measurement precision of our protocol can approach the corresponding QCRB, even for $t\lesssim t_{\rm opt}$.
@@
Furthermore, an important capability of the Floquet-engineered anti-TAT-and-turn readout is its robustness against detection noise. In Fig. 3 (b), we show that our protocol is much more robust than using parity measurement. Here, we assume the detector possessing Gaussian noise and so that the measurement outcome is described by $\langle{\hat{J}}_{\vec{m}}\rangle_{\sigma}=\sum_{m=-N/2}^{N/2}P_{m}(\phi|\sigma)m$. Here $P_{m}(\phi|\sigma)=\sum_{n=-N/2}^{N/2}C_{n}e^{-(m-n)^{2}/2\sigma^{2}}P_{n}(\phi)$ represents the probability distribution under detection noise, $C_{n}=1/\sum_{m=-N/2}^{N/2}\mathrm{e}^{-(m-n)^{2}/2\sigma^{2}}$ is the normalization factor and $P_{n}(\phi)$ denotes the ideal probability distribution. Especially for $\sigma\geq 1$, the parity measurement becomes invalid and our protocol can still attain metrological gain beyond the SQL. Our Floquet-engineered anti-TAT-and-turn can be achieved by adjusting the periodic coupling without changing the interaction, which is feasible for practical entanglement-enhanced sensing.
@@
Conclusions.— We use the continuous Floquet engineering with periodically-modulated Rabi frequency under appropriate detuning to achieve an effective TAT-and-turn dynamics. With semi-classical and quantum approaches, we analytically find the preparation timescale at large particle number is significantly shorter compared to the OAT, and even faster than using artificial three-body interactions [93]. Moreover, applying another Floquet-engineered TAT-and-turn dynamics, one can realize the time-reversal readout without flipping the sign of the nonlinear interaction. Our work provides an alternative way for fast generation and detection of large-particle-number non-Gaussian entangled states, which can be created in various systems [96, 97, 98, 99, 100, 101, 102, 103, 104] and has promising applications in quantum metrology and quantum information science.
@@ Acknowledgements.
The authors would like to thank Dr. Wenjie Liu for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12025509) and the National Key Research and Development Program of China (Grant No. 2022YFA1404104).
@@
By periodically modulating an external field coupled to an ensemble of identical Bose condensed two-level atoms, the Floquet Engineered (FE) Hamiltonian $\hat{H}_{\rm FE}$ is written as
@@
where $\Omega(t)=\Omega_{0}\cos(\omega t)$, and $\hat{J}_{\alpha}=\hat{J}_{x}\cos\alpha+\hat{J}_{y}\sin\alpha$. To perform the analytic calculation, we first transform from the Schrödinger picture to the interaction picture, rewriting $\hat{H}_{\rm FE}$ as a constant term, $\hat{H}_{0}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}$, plus a time-dependent term, $\hat{V}(t)=\Omega_{0}\cos(\omega t)\hat{J}_{\alpha}$. Applying the Baker-Campbell-Hausdorff (BCH) formula, we obtain $e^{i\gamma\hat{J}_{\alpha}}\hat{J}_{z}e^{-i\gamma\hat{J}_{\alpha}}=\cos\gamma\hat{J}_{z}+\sin\gamma\hat{J}_{\beta}$ with $\beta=\alpha+\pi/2$. In the interaction picture, we derive
@@
According to the Euler formula $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and the Jacobi-Anger expansion variant $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}\left[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}\right]$ , where $\mathcal{{J}}_{n}(z)$ is the $n$-th Bessel function of the first kind, we perform the Fourier decomposition and the Hamiltonian can be expressed as
@@
with $\hat{J}_{1,2}=\hat{J}_{z}\pm i\hat{J}_{\beta}$, $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ and $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$. When $n$ is odd (denoted by $p$), we have $H_{-p}^{I}=-H_{p}^{I}$, so that the commutation relations between Hamiltonians are
@@
where $[\hat{P},\hat{Q}]=\hat{P}\hat{Q}-\hat{Q}\hat{P}$ and $\{\hat{P},\hat{Q}\}=\hat{P}\hat{Q}+\hat{Q}\hat{P}$ are commutation and anti-commutation, respectively. In addition, when $n$ is even (denoted by $q$), we have $H_{-q}^{I}=H_{q}^{I}$, the commutation relations become
@@
It indicates that the $q$(even)-order terms make no contribution to the effective Floquet Hamiltonian. Further, in order to keep only $\hat{H}_{0}^{I}$ in the Hamiltonian (16), corresponding to drop the time-dependent terms in the Hamiltonian (11), we require $\omega$ to be sufficiently large $(\omega\gg\chi,\delta)$. Then the Hamiltonian becomes
@@
which reveals that the external driving field leads to the twisting effect along both $z$ and $\beta$ directions. This can be interpreted intuitively as the rotation of spins perpendicular to the axis of OAT ($z$ axis) diverted from the twisting axis. Rewriting the Hamiltonian by adding a constant $-\frac{\chi}{2}(1+L_{0})\hat{J}^{2}$ ($\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}$ is conserved during the dynamics), we obtain a mixture of an OAT Hamiltonian and a TAT Hamiltonian as
@@
When $L_{0}=-1/3$, corresponding to ${\Omega_{0}}/{\omega}\simeq 1.6262$, the Hamiltonian becomes effective TAT-and-turn Hamiltonian
@@
where the effective nonlinear interaction strength $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning $\delta_{\mathrm{eff}}=\delta{K}_{0}\simeq 0.404\delta$. Further, by adjusting the frequency of Rabi frequency and the phase of external coupling field ($\alpha\mapsto\alpha+\pi/2$ and $\delta\mapsto-\delta$), one can obtain the anti-TAT-and-turn Hamiltonian
@@
which can be applied for time-reversal interaction-based readout without flipping the nonlinear interaction.
II II. The maximum quantum Fisher information
In this section, we explain the method of diagonalizing the matrix to obtain the maximum QFI. In quantum metrology, the precision of estimating a parameter $\phi$ is bounded by the Quantum Cramér-Rao Bound (QCRB), given by
$\Delta\phi\geq\frac{1}{\sqrt{\nu F_{Q}}},$ (21)
where $\nu$ is the number of identical experiments, $F_{Q}$ is quantum Fisher information (QFI). For pure states, the QFI is written as
$F_{Q}=4(\Delta\hat{J}_{\vec{n}})^{2}=4(\langle\hat{J}^{2}_{\vec{n}}\rangle-\langle\hat{J}_{\vec{n}}\rangle^{2}),$ (22)
where $\hat{J}_{\vec{n}}$ is the generator for sensing with $e^{-i\hat{J}_{\vec{n}}\phi}$. For a given state, the optimal QFI $F_{Q}$ is determined by the optimal signal encoding direction. For obtaining the optimal signal encoding direction, one can decompose the $\hat{J}_{\vec{n}}=n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z}$ with $n_{x}^{2}+n_{y}^{2}+n_{z}^{2}=1$. Therefore the QFI is written as
$F_{Q}=$ $4\langle(n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z})^{2}\rangle-4\langle n_{x}\hat{J}_{x}+n_{y}\hat{J}_{y}+n_{z}\hat{J}_{z}\rangle^{2}$ (23)
$=$ $n_{x}n_{x}(\langle 2\{\hat{J}_{x},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{x}\rangle)+n_{x}n_{y}(2\langle\{\hat{J}_{x},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{y}\rangle)+n_{x}n_{z}(2\langle\{\hat{J}_{x},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{x}\rangle\langle\hat{J}_{z}\rangle)$
$+n_{y}n_{x}(2\langle\{\hat{J}_{y},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{x}\rangle)+n_{y}n_{y}(2\langle\{\hat{J}_{y},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{y}\rangle)+n_{y}n_{z}(2\langle\{\hat{J}_{y},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{y}\rangle\langle\hat{J}_{z}\rangle)$
$+n_{z}n_{x}(2\langle\{\hat{J}_{z},\hat{J}_{x}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{x}\rangle)+n_{z}n_{y}(2\langle\{\hat{J}_{z},\hat{J}_{y}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{y}\rangle)+n_{z}n_{z}(2\langle\{\hat{J}_{z},\hat{J}_{z}\}\rangle-4\langle\hat{J}_{z}\rangle\langle\hat{J}_{z}\rangle).$
To find the maximum QFI, one needs to optimize over all possible directions $\vec{n}$. This optimization problem can be recast in terms of a matrix diagonalization problem, where the QFI can be represented in matrix form:
$F_{Q}=\begin{pmatrix}n_{x}&n_{y}&n_{z}\\ \end{pmatrix}\begin{pmatrix}\mathcal{F}_{xx}&\mathcal{F}_{xy}&\mathcal{F}_{xz}\\ \mathcal{F}_{yx}&\mathcal{F}_{yy}&\mathcal{F}_{yz}\\ \mathcal{F}_{zx}&\mathcal{F}_{zy}&\mathcal{F}_{zz}\\ \end{pmatrix}\begin{pmatrix}n_{x}\\ n_{y}\\ n_{z}\end{pmatrix},$ (24)
with the Quantum Fisher Information Matrix (QFIM) $\mathcal{F}$ defined as [91]:
$\mathcal{F}=\begin{pmatrix}\mathcal{F}_{xx}&\mathcal{F}_{xy}&\mathcal{F}_{xz}\\ \mathcal{F}_{yx}&\mathcal{F}_{yy}&\mathcal{F}_{yz}\\ \mathcal{F}_{zx}&\mathcal{F}_{zy}&\mathcal{F}_{zz}\\ \end{pmatrix},$ (25)
where the elements are given by
$\mathcal{F}_{\mu\nu}=2\langle\{\hat{J}_{\mu},\hat{J}_{\nu}\}\rangle-4\langle\hat{J}_{\mu}\rangle\langle\hat{J}_{\nu}\rangle,$ (26)
with $\mu,\nu=x,y,z$. Therefore, the maximum QFI $F_{Q}^{\mathrm{max}}$ is the largest eigenvalue of QFIM,
$\mathcal{F}\vec{n}_{\mathrm{max}}=\lambda_{\mathrm{max}}\vec{n}_{\mathrm{max}}=F_{Q}^{\mathrm{max}}\vec{n}_{\mathrm{max}},$ (27)
where the optimal signal encoding direction vector $\vec{n}_{\mathrm{max}}=(n_{x}^{\mathrm{max}},n_{y}^{\mathrm{max}},n_{z}^{\mathrm{max}})^{\mathrm{T}}$, and the responding optimal generator is $\hat{J}_{\vec{n}_{\mathrm{max}}}=n_{x}^{\mathrm{max}}\hat{J}_{x}+n_{y}^{\mathrm{max}}\hat{J}_{y}+n_{z}^{\mathrm{max}}\hat{J}_{z}$.
III III. Semi-classical treatment of TAT-and-turn dynamics
We use the semi-classical treatment, introduced in Ref. [92], to analyze the metrological properties of TAT-and-turn dynamics. Utilizing the commutations $[\hat{J}_{\alpha},\hat{J}_{\beta}]=i\epsilon_{\alpha\beta z}\hat{J}_{z}$ or its variants, we can describe the time-evolution of Hamiltonian (19) by the Heisenberg equations of motion, $d\hat{J}_{\kappa}/dt=i[\hat{H},\hat{J}_{\kappa}]$ where $\kappa=\alpha,\beta,z$ are the components of the collective spin. They are
@@
In the thermodynamic limit, $J=N/2\to\infty$, above Eqs.(28) lead to the phase-space flow of the classical variables $\mathbf{R}=(A,B,Z)=$ $(\langle\hat{{J}}_{\alpha}\rangle,\langle\hat{{J}}_{\beta}\rangle,\langle\hat{{J}}_{z}\rangle)/J$ with $A^{2}+B^{2}+Z^{2}=1$, which after neglecting correlations $\langle\hat{P}\hat{Q}\rangle=\langle\hat{P}\rangle\langle\hat{Q}\rangle$, is given by
@@
Likewise, the classical phase-space trajectories of nonlinear or linear modulation are plotted in Fig.1 (c) of the main text.
The direction of the initial spin coherent state of the non-adiabatic evolution of the Hamiltonian (19) is determined by fixed points of a phase-space flow with a trivial evolution. By solving $d\mathbf{R}/dt=0$, the phase-space flow in Eqs. (29) has six different fixed points depending on the system parameters (we consider the conditions about $\chi\neq 0$, and $|\delta_{\mathrm{eff}}/(N\chi_{\mathrm{eff}})|<1$) are given by
@@
For the fixed points, the stability analysis comes from the diagonalization of the Jacobi matrix, ${M}[\mathbf{R}]=(\partial/\partial\mathbf{R})(d\mathbf{R}/dt)$. In Eqs. (29), this matrix of the phase-space flow is given by
@@
The analysis shows that two fixed points $(A,B,Z)=(0,0,\pm 1)$ that always pointing toward the ±$\pm z$-axis are saddle points and the other four points are stable points. The Jacobi matrix of the saddle points is
@@
Figure 4: The maximum QFI $F_{Q}^{\mathrm{max}}$ obtained via TAT-and-turn dynamics varying in short evolution time with different $\delta$ for particle number $N=100$. The white dashed line represents the maximum varying in time, and indicates that the QFI has the fastest growth rate within short evolution time when $\delta_{\mathrm{eff}}=0$.
In general, the separatrix branches emerge from this saddle point, which yield the classical evolution trajectory of the Hamiltonian. By evaluating the eigenvalues of the Jacobi matrix in Eq. (32) at this saddle, one obtains its local Lyapunov exponent
@@
It dictates the exponential rate at which points move away from the saddle point, i.e., the optimal entanglement generation parameters at the initial moment. This rate reaches maximum when $\delta_{\mathrm{eff}}=0$, which is equivalent to the TAT Hamiltonian and verified to generate metrologically useful states within a short time in Fig. 4, and the corresponding Lyapunov exponent is given by $\Lambda_{\mathrm{sd}}^{\mathrm{TAT}}=N\chi_{\mathrm{eff}}$.
Besides, we can obtain the energy density of the TAT-and-turn Hamiltonian by computing $\langle\hat{H}_{\mathrm{TATNT}}\rangle/J$ in the thermodynamic limit. After neglecting correlations as before and in terms of the classical variables $\mathbf{R}=(A,B,Z)$, this energy density in classical phase space becomes
@@
Since the conservation of energy guarantees that points on the separatrix have the same energy as the one of the saddle, the separatrix equation of two saddles can be constructed by $E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=E(0,0,\pm 1;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\pm\delta_{\mathrm{eff}}$. Due to $A^{2}+B^{2}+Z^{2}=1$, the above formula can be rewritten as
@@
The relationship between classical space components is obtained as

### v2

@@ Abstract
Interaction-based readout without flipping the sign of nonlinear interaction.— It is crucial to approach the ultimate precision bound (i.e. the QCRB) set by the QFI. For a tiny estimated phase, time-reversal readout may approach the QCRB without single-particle resolved detection [56, 57, 58, 59, 55, 60, 61, 62, 63, 80, 81, 64, 65, 66]. Here, we find that the optimal measurements can be achieved via anti-TAT-and-turn, regarded as the time-reversal process of TAT-and-turn. More importantly, the anti-TAT-and-turn can be realized by Floquet engineering without the need of flipping the sign of nonlinear interaction $\chi$.
@@
With the generated GHZ-like state, one can encode the estimated phase $\phi$ along the best sensing direction by $\hat{R}_{\vec{n}_{\rm max}}(\phi)=e^{-i\phi\hat{J}_{\vec{n}_{\rm max}}}$. Here, $\phi$ is accumulated during free evolution and depends on the physical quantity to be measured. The physical quantity can be the frequency detuning of a reference laser driving the atoms in atomic clocks or the magnetic field responsible for Zeeman splitting of the atomic levels in magnetometers. Based on the Hamiltonian (11), by adjusting the detuning $\delta\mapsto-\delta$ and the initial phase of external coupling field $\alpha=0\mapsto\alpha=\pi/2$, one can obtain an effective anti-TAT-and-turn Hamiltonian $\hat{H}_{\rm{TATNT}}^{\alpha=\pi/2,-\delta}=-\hat{H}_{\rm{TATNT}}^{\alpha=0,\delta}$. Ideally, the state before measurement can be written as
@@
where $\lvert\psi_{i}\rangle=\lvert\uparrow\rangle^{\otimes N}$, $\hat{H}_{\rm TATNT}^{\alpha=0,\delta}=\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})+\delta_{\rm eff}\hat{J}_{z}$ and $\hat{H}_{\rm{TATNT}}^{\alpha=\pi/2,-\delta}=-\chi_{\rm eff}(\hat{J}_{y}^{2}-\hat{J}_{x}^{2})-\delta_{\rm eff}\hat{J}_{z}$. Furthermore, to find out the optimal measurement, one can optimize the sensitivity over all possible directions through singular value decomposition
@@
where $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ with two $3\times 3$ matrices $\mathcal{M}$ and $\mathcal{Q}$, whose elements are given as $\mathcal{M}_{\mu\nu}=i\langle[\hat{U}^{\dagger}_{1}\hat{J}_{\mu}\hat{U}_{1},\hat{\tilde{J}}_{\nu}]\rangle_{i}$ and $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle_{i}/2-\langle\hat{\tilde{J}}_{\mu}\rangle_{i}\langle\hat{\tilde{J}}_{\nu}\rangle_{i}$ for $\mu,\nu=x,y,z$ with $\hat{\tilde{J}}_{\mu(\nu)}=\hat{U}^{\dagger}_{1}\hat{R}_{\vec{n}_{\rm max}}^{\dagger}(\phi)\hat{U}^{\dagger}_{2}\hat{J}_{\mu(\nu)}\hat{U}_{2}\hat{R}_{\vec{n}_{\rm max}}(\phi)\hat{U}_{1}$ [98, 99]. The optimal measurement direction can then be given as $\vec{m}_{\rm max}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with the normalization constant $\mathcal{N}_{c}$ determined by the Cauchy-Schwarz inequality [53]. Therefore, one can easily rotate the final state and measure the half-population difference $\hat{J}_{z}$ to infer $\phi$. In particular, the above calculations are valid for an arbitrary phase $|\phi|\geq 0$, which generalizes the result only for the zero phase $\phi=0$ [73]. As shown in Fig. 3 (a), the nonlinear readout without flipping the sign of nonlinear interaction for FE-TATNT achieves excellent measurement precision approaching the corresponding QCRB.
@@
Furthermore, an important capability of the Floquet-engineered anti-TAT-and-turn readout is its robustness against detection noise. Here, we assume the detector possessing Gaussian noise and the probability distribution under detection noise is described by $P_{m}(\phi|\sigma)=\sum_{n=-N/2}^{N/2}C_{n}e^{-(m-n)^{2}/2\sigma^{2}}P_{n}(\phi)$ with the normalization $C_{n}=1/\sum_{m=-N/2}^{N/2}\mathrm{e}^{-(m-n)^{2}/2\sigma^{2}}$ and the ideal probability distribution $P_{n}(\phi)=|\langle n\lvert\psi_{f}(\phi)\rangle|^{2}$ [58]. Under detection noise, the outcome is expressed as $\langle{\hat{J}}_{z}\rangle_{\sigma}=\sum_{m=-N/2}^{N/2}mP_{m}(\phi|\sigma)$ for population measurement or $\langle\Pi\rangle_{\sigma}=\sum_{m=-N/2}^{N/2}(-1)^{m}P_{m}(\phi|\sigma)$ for parity measurement [61]. As shown in Fig. 3 (b), for $\sigma\geq 1$, while the parity measurement becomes invalid, our protocol can still achieve a metrological gain beyond SQL, showing that the robustness of our protocol is much better than that of using the parity measurement. Our Floquet-engineered anti-TAT-and-turn can be achieved by adjusting the periodic coupling without changing the interaction, which is feasible for realistic entanglement-enhanced sensing.
@@
Conclusions.— We use continuous Floquet engineering with periodically modulated Rabi frequency under appropriate detuning to achieve an effective TAT-and-turn dynamics. With semi-classical and quantum approaches, we analytically find the preparation timescale at large particle number is significantly shorter compared to the OAT, and even faster than using artificial three-body interactions, which is advantageous in mitigating decoherence. Moreover, applying another Floquet-engineered TAT-and-turn dynamics, one can realize the time-reversal readout without flipping the sign of the nonlinear interaction. Our work provides an alternative way for fast generation and detection of large-particle-number GHZ-like states, which can be created in various systems [84, 85, 86, 87, 88, 89, 90, 91, 92, 93] and has promising applications in quantum metrology and quantum information science.
@@ Acknowledgements.
The authors thank Dr. Wenjie Liu for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grants No. 12025509 and No. 92476201), the National Key Research and Development Program of China (Grant No. 2022YFA1404104), and the Guangdong Provincial Quantum Science Strategic Initiative (GDZX2305006 and GDZX2405002).
@@
By periodically modulating an external field that couples to an ensemble of identical Bose-condensed two-level atoms, the Floquet-engineered Hamiltonian can be expressed as:
@@
where $\Omega(t)=\Omega_{0}\cos(\omega t)$, and $\hat{J}_{\alpha}=\hat{J}_{x}\cos\alpha+\hat{J}_{y}\sin\alpha$. To facilitate the analytical calculation, we first transform from the Schrödinger picture to the interaction picture, where $\hat{H}_{\rm FE}$ is decomposed into a time-independent term, $\hat{H}_{0}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}$, and a time-dependent term, $\hat{V}(t)=\Omega_{0}\cos(\omega t)\hat{J}_{\alpha}$. Applying the Baker-Campbell-Hausdorff formula, we obtain $e^{i\gamma\hat{J}_{\alpha}}\hat{J}_{z}e^{-i\gamma\hat{J}_{\alpha}}=\cos\gamma\hat{J}_{z}+\sin\gamma\hat{J}_{\beta}$ where $\beta=\alpha+\pi/2$. In the interaction picture, we derive
@@
Using the Euler formula $e^{i\gamma}=\cos{\gamma}+i\sin{\gamma}$ and the Jacobi-Anger expansion variant $e^{iz\sin\theta}=\mathcal{{J}}_{0}(|z|)+\sum_{n=1}^{\infty}\left[\mathcal{{J}}_{n}(z)e^{in\theta}+(-1)^{n}\mathcal{{J}}_{n}(z)e^{-in\theta}\right]$ , where $\mathcal{{J}}_{n}(z)$ is the $n$-th Bessel function of the first kind, we perform the Fourier decomposition and the Hamiltonian can be expressed as
@@
with $\hat{J}_{1,2}=\hat{J}_{z}\pm i\hat{J}_{\beta}$, $L_{n}=\mathcal{{J}}_{n}\left(2\Omega_{0}/\omega\right)$ , $K_{n}=\mathcal{{J}}_{n}\left(\Omega_{0}/\omega\right)$ and $n=1,2,3\cdots$. When $n$ is odd (denoted by $p$), we have $H_{-p}^{I}=-H_{p}^{I}$, so the commutation relations between Hamiltonians are
@@
where $[\hat{P},\hat{Q}]=\hat{P}\hat{Q}-\hat{Q}\hat{P}$ and $\{\hat{P},\hat{Q}\}=\hat{P}\hat{Q}+\hat{Q}\hat{P}$ are commutation and anti-commutation, respectively. For even $n$ (denoted by $q$), we have $H_{-q}^{I}=H_{q}^{I}$, and the commutation relations become
@@
It indicates that the $q$(even)-order terms do not contribute to the effective Floquet Hamiltonian. In order to retain only $\hat{H}_{0}^{I}$ in the Hamiltonian (18), corresponding to drop the time-dependent terms in the Hamiltonian (13), we require $\omega$ to be sufficiently large $(\omega\gg N\chi)$. Under this condition, the Hamiltonian simplifies to
@@
which reveals that the external driving field generates twisting along both the $z$ and $\beta$ directions. Assuming that the $\hat{J}^{2}=\hat{J}_{\alpha}^{2}+\hat{J}_{\beta}^{2}+\hat{J}_{z}^{2}$ is conserved during the dynamical evolution, we derive a Hamiltonian that combines features of both OAT and TAT Hamiltonians by incorporating an additional term $-\frac{\chi}{2}(1+L_{0})\hat{J}^{2}$. The resulting Hamiltonian can be expressed as:
@@
When $L_{0}=-1/3$, corresponding to ${\Omega_{0}}/{\omega}\simeq 1.6262$, the Hamiltonian becomes an effective TAT-and-turn Hamiltonian:
@@
where the effective nonlinear interaction strength is $\chi_{\mathrm{eff}}=\chi/3$ and the effective detuning is $\delta_{\mathrm{eff}}=\delta{K}_{0}\simeq 0.4404\delta$. Furthermore, by adjusting the initial phase of external coupling field and flipping the detuning ($\alpha\mapsto\alpha+\pi/2$ and $\delta\mapsto-\delta$), one can obtain the anti-TAT-and-turn Hamiltonian
@@
which can be applied for time-reversal interaction-based readout without flipping the sign of nonlinear interaction.
II II. Semi-classical treatment of TAT-and-turn dynamics
We use the semi-classical treatment, as introduced in Ref. [96], to analyze the metrological properties of TAT-and-turn dynamics. Utilizing the commutation relations $[\hat{J}_{\alpha},\hat{J}_{\beta}]=i\epsilon_{\alpha\beta z}\hat{J}_{z}$, we can describe the time-evolution of Hamiltonian (21) using the Heisenberg equations of motion, $d\hat{J}_{\kappa}/dt=i[\hat{H},\hat{J}_{\kappa}]$, where $\kappa=\alpha,\beta,z$ are the components of the collective spin. The Heisenberg equations of motion are given by
@@
In the thermodynamic limit, $J=N/2\to\infty$, above Eq.(23) leads to the phase-space flow of the classical variables $\mathbf{R}=(A,B,Z)=$ $(\langle\hat{{J}}_{\alpha}\rangle,\langle\hat{{J}}_{\beta}\rangle,\langle\hat{{J}}_{z}\rangle)/J$ with $A^{2}+B^{2}+Z^{2}=1$. After neglecting correlations $\langle\hat{P}\hat{Q}\rangle=\langle\hat{P}\rangle\langle\hat{Q}\rangle$, the equations are given by
@@
Likewise, the classical phase-space trajectories of nonlinear OAT or linear modulation can be obtained using the above method and are plotted in Fig.1 (c) of the main text.
The direction of the initial spin coherent state in the non-adiabatic evolution of the Hamiltonian (21) is determined by the fixed points of a phase-space flow with a trivial evolution. By solving $d\mathbf{R}/dt=0$, the phase-space flow in Eq. (24) has six different fixed points depending on the system parameters (we consider the conditions $\chi\neq 0$, and $|\delta_{\mathrm{eff}}/(N\chi_{\mathrm{eff}})|<1$), which are given by
@@
For the fixed points, the stability analysis is carried out by diagonalizing of the Jacobi matrix, ${M}[\mathbf{R}]=(\partial/\partial\mathbf{R})(d\mathbf{R}/dt)$. In Eq. (24), the matrix of the phase-space flow is given by
@@
The analysis shows that two fixed points $(A,B,Z)=(0,0,\pm 1)$ which always point along the ±$\pm z$-axis are saddle points, while the other four points are stable. The Jacobi matrix of the saddle points are
@@
In general, the separatrix branches emerge from this saddle point, which yield the classical evolution trajectory of the Hamiltonian. By evaluating the eigenvalues of the Jacobi matrix in Eq. (27) at this saddle, one obtains its local Lyapunov exponent:
@@
It dictates the exponential rate at which points move away from the saddle point, i.e., the optimal entanglement generation parameters at the initial moment. This rate reaches its maximum when $\delta_{\mathrm{eff}}=0$, which corresponds to the TAT Hamiltonian and the corresponding Lyapunov exponent is given by $\Lambda_{\mathrm{sd}}^{\mathrm{TAT}}=N\chi_{\mathrm{eff}}$.
Besides, we can obtain the energy density of the TAT-and-turn Hamiltonian by computing $\langle\hat{H}_{\mathrm{TATNT}}\rangle/J$ in the thermodynamic limit. After neglecting correlations as before, and in terms of the classical variables $\mathbf{R}=(A,B,Z)$, the energy density in classical phase space becomes
@@
Since the conservation of energy ensures that points on the separatrix have the same energy, the separatrix equation of the two saddles can be constructed as $E(A,B,Z;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=E(0,0,\pm 1;\delta_{\mathrm{eff}},\chi_{\mathrm{eff}})=\pm\delta_{\mathrm{eff}}$. Due to the constraint $A^{2}+B^{2}+Z^{2}=1$, the above formula can be rewritten as
@@
The relationships between classical space components are obtained as